OpenStudy (anonymous):

Let y=(7/x^3)(√196x^6 + 364x^2 + 169/(x^2)), x>0. Find dy/dx and simplify completely. Hint:Before finding derivative, simplify y!

OpenStudy (anonymous):

\[y = \frac{ 7 }{ x^3}\sqrt{196x^6 + 364x^2 + \frac{ 169 }{ x^2 }}\]